### v1

# Towards Characterizing Subjectivity of Individuals through Modeling Value Conflicts and Trade-offs

Abstract
Large Language Models (LLMs) not only have solved complex reasoning problems but also exhibit remarkable performance in tasks that require subjective decision making. Existing studies suggest that LLM generations can be subjectively grounded to some extent, yet exploring whether LLMs can account for individual-level subjectivity has not been sufficiently studied. In this paper, we characterize subjectivity of individuals on social media and infer their moral judgments using LLMs. We propose a framework, SolAr (Subjective Ground with Value Abstraction), that observes value conflicts and trade-offs in the user-generated texts to better represent subjective ground of individuals. Empirical results show that our framework improves overall inference results as well as performance on controversial situations. Additionally, we qualitatively show that SolAr provides explanations about individuals’ value preferences, which can further account for their judgments.
Towards Characterizing Subjectivity of Individuals through
Modeling Value Conflicts and Trade-offs
Younghun Lee  and Dan Goldwasser Department of Computer Science Purdue University West Lafayette, IN, USA {younghun, dgoldwas}@purdue.edu
1 Introduction
For the last few years, Large Language Models (LLMs) have shifted the paradigm of solving NLP problems to autoregressive language generation and achieved human-like performance in many downstream tasks (Raffel et al., 2020; Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022; Yu et al., 2022; He et al., 2024). Not only have LLMs solved objective problems that require complex reasoning skills, such as STEM-related questions (Imani et al., 2023; Wang et al., 2024c; Abbasiantaeb et al., 2024), they also exhibit remarkable performance in subjective decision-making processes, such as bias detection (Hartvigsen et al., 2022), representing opinions in the survey (Santurkar et al., 2023), generating model evaluation (Perez et al., 2022), social context grounding (Pujari et al., 2024), etc.
Recent studies explore whether LLMs can generate perspectives and reasoning that align well with a specific persona or demographic information (Durmus et al., 2023; Nie et al., 2024; Zheng et al., 2024b). The results of these studies show that it is possible, to an extent, to ground LLM generations with these traits, however, LLMs tend to rely on superficial facts and assumptions about the roles rather than apply a deeper understanding.
Our goal in this paper is to study a different aspect of subjectivity, focusing on the individual-level (rather than generalizing over demographic traits), which has not been sufficiently studied yet. As personalized AI becomes more widely used (McClain, 2024), understanding whether LLMs can be utilized to characterize individual subjectivity becomes more important. Analyzing individual-level subjectivity using LLMs faces two main challenges. The first challenge is guiding LLM generation to be consistent with a specific subjective view. Existing methods for capturing subjectivity at the level of a generalized persona, role, or demographic information show that it is not trivial to steer LLMs to follow certain aspects. Second, even if an optimal approach for grounding subjective aspects in LLMs generations existed, expressing these aspects at the level of an individual is not straightforward (i.e., two instances of the same persona could still have different subjective preferences). Conceptualizing and operationalizing subjectivity at this level poses a second challenge.
Figure 1: Inference process of our framework, SolAr. When an input situation is given, SolAr identifies conflicting values in the story, and retrieve the most relevant value trade-off history from the redditor’s past comments. The retrieved subjective ground is then added to the prompt to infer each redditor’s likely judgments to the situation.
In this paper, our objective is to characterize subjectivity of individuals with LLMs by analyzing users’ behaviors in a Reddit community, r/AmITheAsshole. In this community, original posters write about the situations where they have conflicts with others and ask whether their behaviors are acceptable, and other redditors leave comments with their judgments. Figure 1 shows an example of a post and judgments from different redditors. The research hypothesis is that redditors’ subjective ground, principles that play a fundamental role in human moral judgments (Neuhouser, 1990), can be represented by decomposing their past behaviors; we use redditors’ past comments in the community to determine their likely judgments on unseen situations.
Value pluralism suggests that there are multiple values that may be equally correct, yet in conflict with one another (Crowder, 1998; Galston, 2002). Psychological studies adopt this idea and argue that the human cognitive system makes novel judgments by making trade-offs between conflicting values when it encounters situations with colliding moral intuitions (Fiske and Tetlock, 1997; Guzmán et al., 2022). In this paper, we propose a framework, SolAr (Subjective Ground with Value Abstraction), that observes trade-offs between conflicting values in the user-generated texts, identifies the most relevant value conflicts with respect to the input situation, and infer the redditors’ likely judgments using off-the-shelf LLMs. Our framework aims to tackle the two challenges of characterizing individual-level subjectivity described above; it conceptualizes subjectivity of individuals with trade-offs among abstract values and guides LLM generations with Retrieval-Augmented Generation (RAG). Empirical results suggest that SolAr distinguishes redditors’ subjective preferences and further provides explanations of their value trade-offs. Our proposed framework also improves the overall performance of downstream tasks and better guides LLMs with the prompted subjective ground when it is tested on more controversial situations.
Key Contributions: To the best of our knowledge, this is the first attempt to explore whether off-the-shelf LLMs can be effectively used to account for subjectivity of individuals in the online community. With value abstraction, we highlight that redditors show distinct subjectivity patterns, which makes this research different from social commonsense or opinion mining. Additionally, we propose a novel framework that encompasses a value trade-off system and performs better in downstream tasks as well as grounding LLM generations.
2 Problem Formulation
In this section, we describe four major elements that formulate modules, learning processes, and inference tasks.
Situation Situation refers to the text description of what has happened in the real world. We aim to focus on situations that portray conflicts so that one’s point of view can be projected in diverse ways. “Asking my girlfriend to stay out of apartment for poker night” is an example of situations.
Individual This research aims to analyze different individuals’ reactions and responses on various social situations. Unlike opinion mining or social commonsense studies, where perspectives and rationales are represented as an aggregate of a large number of humans, our main focus is to observe distinct patterns that characterize each individual and understand the rationales behind their decision-making processes.
Subjective Ground Subjective Ground refers to principles or maxims that steer individuals’ moral judgments or perspectives. “One should put their significant other’s needs as top priority.” is a subjective ground item that is relevant to the example situation above. Every individual has a distinct subjective ground and is built from various aspects such as their past experience, demographics, personality, etc. (Eysenck and Eysenck, 1975; Schwaba et al., 2023; Schoeller et al., 2024). One of our research hypotheses is that the subjective ground of individuals can be inferred by observing their past behaviors. Throughout this research, we aim to validate this hypothesis by modeling the subjective ground with individuals’ comment history on social media.
Subjectivity Abstraction One of the limitations of the aforementioned assumption is that it works in an ideal setting in which individuals’ past behaviors can be observed across an extremely large range of social situations—an impractical scenario in real-world contexts. Thus in reality, when inferring individuals’ likely judgments on unseen situations, it is necessary to formulate some hypotheses based on observable subjective ground. Subjectivity Abstraction refers to a process that produces a high-level representation of subjective ground that can be generalized to a broader spectrum of situations. In this paper, we explore several approaches for constructing subjectivity abstraction.
3 Task Description
We analyze subjective perspectives and judgments of individuals posted on a Reddit community, r/AmITheAsshole. In this community, original posters (OP) describe situations in which they have conflicts with others and ask if they are at fault. Other individuals (i.e., redditors) then leave comments and judge the acceptability of the OP’s behaviors in the situation. We aim to analyze the subjectivity of individuals who leave comments and judgments about the situations.
There are a couple of benefits to using this community in analyzing individual-level subjectivity with language models. First, the situations described in this community are mostly about everyday events that are generic (e.g. “not attending a friend’s wedding”) rather than related to specific world events (e.g. “commenting on the new executive orders from the president”); thus, the language models can have a better understanding of the situations without having a knowledge gap. Another benefit is that the redditors’ subjective judgments are coded in a discrete fashion, which makes language model predictions more objective compared to open-ended analysis of subjectivity.
3.1 Crawling from r/AmITheAsshole
We crawl all posts in the r/AmITheAsshole community from November 2014 to June 2023 and identify the top 8 redditors who commented the most to ensure a sufficient amount of data for each individual. To ensure that the comments solely mention the situations, we consider top-level comments without including any threaded comments. We then filter out the posts for which the top 8 users did not leave their judgments.
In the r/AmITheAsshole community, redditors leave the judgments in their comments with pre-coded words; YTA (You’re The Asshole), YWBTA (You Would Be The Asshole), NTA (Not The Asshole), YWNBTA (You Would Not Be The Asshole), ESH (Everyone Sucks Here), NAH (No Assholes Here), and INFO (Not Enough Info). For simplicity, we group NTA, NAH, and YWNBTA as ‘acceptable’, and YTA, ESH, and YWBTA as ‘unacceptable’. INFO is discarded as it does not convey subjectivity. Detailed statistics of the crawled dataset are described in Table 1.
Crawled r/AmITheAsshole
# of total instances 18,669
# of unique situations 17,432
Max / Min # of instances per redditor 2,857 / 1,782
# Accept. / Unaccept. labels (overall) 12,279 / 6,390
# Accept. / Unaccept. labels (skewed) 3,251 / 269
Table 1: Statistics of the dataset. The dataset takes the most active 8 redditors into consideration, keeping the instances that contain coded judgments in the comments. Label distributions are described in two ways; by combining all redditors (overall), and by combining two redditors who showed the most skewed judgment patterns (skewed).
3.2 Moral Value Annotation
As discussed in 2, we produce a high-level abstraction of redditors’ comments and the situations. For each pair of situation and an individual’s comment, we prompt off-the-shelf LLMs and generate values that are observed from the situation and redditors’ comments. As human values can be defined and captured differently in the same text, we apply several different approaches.
We first analyze and annotate the texts based on the theory of basic human values proposed by Schwartz (1992). The theory suggests ten basic values that could explain how people in different cultures recognize the underlying motivation and goals. One of the advantages of using this framework is that it explains values that align with or conflict against one another, which naturally expands hypotheses of individual subjectivity. For instance, when we observe an individual’s preference of “Openness to Change”, it is implied that the individual is likely to have less preference of “Conservation” or “Tradition”. Detailed explanations of the ten basic human values and how they are annotated are described in Appendix A.1.
As opposed to using a fixed framework for characterizing human values, we use LLMs to generate more open-ended text of the values observed in the texts. In this setup, we apply value trade-off theories. Humans make judgments based on value trade-offs when different values conflict to each other (Fiske and Tetlock, 1997; Leyva, 2019; Guzmán et al., 2022). As situations in the dataset mostly describe conflicts OPs are having, we prompt LLMs to identify all conflicting value pairs in the situation and discover value trade-offs made by each redditor in the comments.
When conflicting values are generated in an open-ended manner by LLMs, the level of abstraction is insufficient; values describe situation-specific details (e.g. “prioritizing girlfriend’s plan to cook together”), rather than general concepts (e.g. “partner’s emotional needs”). To address this, we iteratively cluster similar value representations and discover high-level definitions for these clustered values. The initial clusters are formed using HDBSCAN (McInnes et al., 2017), and we later use LLMs to create additional clusters for values that remain uncategorized in the initial clustering phase, resulting in 111 clusters in total. Detailed processes for value generation and clustering are described in Appendix A.2.
3.3 Learning Problem
Our major focus is to use language models as a reasoning machine to understand the subjective ground of individuals and predict their likely behaviors in unseen instances. More specifically, we formulate a binary classification task where the language models are given with situations and redditors’ most relevant subjective ground, either their past comments or value abstractions, and generate a prediction whether the redditors would judge the OP’s behaviors described in input situations acceptable or not.
4 Model
In this research, we propose a framework, SolAr, that accomplishes the task with Retrieval-Augmented Generations (RAG).
4.1 Subjective Ground Retrieval
In order to infer how the target individual would react to the given input, the most relevant subjective ground needs to be retrieved. We design several heuristics to operationalize the retrieval. First, we could simply select the comments that are left on situations similar to the test situations by computing the pairwise distances between the vector representations of the situations.
While retrieving comments from similar situations, we add another heuristics to explicitly retrieve comments that show different judgment patterns and guarantee that LLMs are prompted with different moral judgments. Suppose that LLMs are prompted to predict whether a redditor judges “not paying for my daughter’s wedding” acceptable or not. In observing the past history, it is possible that this redditor commented to the top 5 most similar situations as “not acceptable”. We then try to find other instances where the redditor says differently. When the redditor judges as “acceptable” to “refusing to contribute to my daughter’s wedding after she cancelled the previous one”, LLMs not only understand the redditor’s general judgment tendencies, but also observe special conditions that would change their usual judgment patterns.
Alternatively to computing the distance between situation representations, we could compute the pairwise distances of the abstract values for all situations. The motivation of using value representations is to retrieve comments from more diverse situations; distance among situation representations yields semantically or topically similar situations only, while value similarity can retrieve situations that are similar not in terms of topics, but in terms of high-level values. Consider an input situation about wedding ceremonies for instance, similarity based on situation representations would result in all wedding related situations, while using representation of abstract values could retrieve situations about social appearance, money, family relationship dynamics, etc.
4.2 Judgment Prediction
Given an input and a set of subjective ground provided by the retrieval process, off-the-shelf LLMs predict the likely moral judgment of a redditor. Prediction language models could be designed as a universal language model; this implies that the difference among individuals mostly comes from their subjective ground, and the output reactions derived from the retrieved subjective ground need to be reasonable in general. In this paper, we use the gpt-4o model (Hurst et al., 2024) to predict the acceptability judgments of individuals.
5 Experiments
In addition to our proposed framework using LLMs for reasoning, we implement trainable language models varying in structures and subjective ground representations for better comparison. Macro F1 score is used as an evaluation metric, as the label distribution is unbalanced.
[…]
The details of the implementation are described in Appendix B, and the codes will be released for future reproduction.
5.1 Baseline Models
[…]
We implement three encoder-only models, DistilBERT (Sanh et al., 2019), RoBERTa (Liu et al., 2019), and DeBERTa-v3 (He et al., 2021), where the input for each model is a text description of the situation and the output is the redditor’s acceptability judgments, 0 or 1.
Moral Judgment Prediction
Model F1-overall F1-skewed
Encoder-only Models
DistilBERT 68.71 52.78
RoBERTa $+$ LoRA 70.48 53.87
DeBERTa-v3 70.94 59.32
Encoder-Decoder Models
BART 68.50 53.55
FLAN-T5 65.44 47.79
Hierarchical Models
BART $+$ DeBERTa 66.67 55.75
FLAN-T5 $+$ DeBERTa 66.86 54.38
gpt-4o as Classifier
Random author’s Comm 72.17 64.53
Same author, random Comm 74.99 73.26
5-shot with Title $+$ Judgment 72.85 66.95
5-shot with Comment 76.78 72.89
5-shot with Diff Judgment 77.43 74.44
SolAr: gpt-4o Clf, Selection w/ Abstraction
Schwartz’s Basic Values 75.81 72.95
[…]
Table 2: Macro F1 scores averaging all redditors (-overall) and two redditors with extremely skewed label distributions (-skewed).
We also fine-tune encoder-decoder language models, mainly BART (Lewis, 2019) and FLAN-T5 (Chung et al., 2024), to learn more about each redditor’s subjectivity. The models get the same input, but their objective is to generate the redditors’ likely reactions (i.e., comments and judgments). While encoder-only models only observe each redditor’s binary judgment patterns with respect to the situations, the models fine-tuned with this approach have access to the texts that are authored by the redditors.
Finally, we implement hierarchical models that combine language generation abilities of encoder-decoder models and classification abilities of encoder-only models. In this setup, encoder-decoder models generate each redditor’s likely reactions without judgments, and a separate encoder-only model learns likely judgments from the generated comments.
Since the number of training instances for each redditor might not be sufficient, we try techniques that help fine-tuning the models with smaller number of instances such as LoRA (Hu et al., 2021). We report the best combinations resulting for each model structure in Table 2.
5.2 Subjective Ground Selection
As described in 4.1, we represent the subjective ground of an individual in four different ways: using comments from similar situations, comments from similar situations while presenting different judgments, abstract values generated from comments using fixed value definitions (i.e., Schwartz’s theory of basic human values), and abstract value trade-offs generated by LLMs. We choose the top five items as few-shot examples for each instance and prompt the gpt-4o model to get the final judgment prediction.
In order to compare the effectiveness of subjective ground retrieval, we prompt the same gpt-4o model with randomized few-shot examples in two ways, random redditors’ comments, and the target redditor’s random comments. Random redditors’ comments are neither relevant to the test situations nor represent subjective preferences of the target redditor, while the target redditor’s random comments provide what the target redditor has commented on other irrelevant situations.
6 Discussion
In this section, we analyze the inference performance of different models and discuss the effectiveness of each model component. In addition, we perform qualitative analyses of each redditor’s subjective ground with respect to annotated abstract values and assess how the proposed framework explains individual-level subjectivity.
6.1 Inference Performance
[…]
We also observe that as the model becomes more complex and parameterized, the inference performance worsens—the F1 scores of encoder-decoder models are worse than encoder-only models, and hierarchical models are even worse than that. This implies that the nature of data scarcity in subjectivity analysis also makes it more difficult to fine-tune a specific individual’s perspectives to characterize their subjectivity.
Selecting the most relevant subjective ground of redditors given a test situation works generally well, answering one of our research questions, “Can off-the-shelf LLMs account for individual’s subjective preference?”. As RAG-based inference does not require training, the performance of F1-skewed becomes more similar to the overall F1 scores.
Among different selection methods, using vector representations of value trade-offs work the best. Mapping comments to a fixed set of value systems suggested by Schwartz (1992) performs worse than using redditors’ comments as is. This is potentially because ten fixed set of values are too abstract to map diverse perspectives of individuals on social situations. Selecting few-shot examples with enforcing different moral judgments perform the second best, implying that it is important to provide more diverse examples of redditors’ subjective preferences in order to better ground LLMs.
Moral Judgment Prediction - Controversial Situations
Model F1-overall F1-skewed
Encoder-only Models
DistilBERT 58.15 42.04
RoBERTa 59.95 49.35
DeBERTa-v3 56.32 56.93
gpt-4o as Classifier
Random Author’s Comm 52.65 56.62
Same author, random Comm 55.40 59.61
5-shot with Title $+$ Judgment 56.45 59.27
5-shot with Comment 55.66 59.27
5-shot with Diff Judgment 57.45 62.30
SolAr: gpt-4o Clf, Selection w/ Abstraction
Schwartz’s Basic Values 55.00 61.39
[…]
Table 3: Macro F1 scores on controversial situations. RAG-based methods, except for our proposed framework, perform worse than fine-tuned encoder-only classifiers.
Figure 2: Heat map of value trade-offs among the top 8 redditors. There are some pairs of conflicting values where the majority of redditors show similar preference (e.g. tradeoff_C), yet in most cases their value trade-offs work differently.
6.2 Performance on Controversial Situations
One of the surprising results from Table 2 is that prompting with random comments still perform reasonably well. We hypothesize that LLMs tend to judge the situations on their own when the few-shot examples no longer provide relevant subjective choices, and their judgment patterns would look similar to humans’ if the situation is less controversial.
[…]
6.3 Do people show diversities in value trade-offs?
A key research hypothesis of this study is that each redditor actively participating in r/AmITheAsshole exhibits distinct subjectivity patterns, making the analysis of individual perspectives a fundamentally different NLP challenge compared to social commonsense reasoning. We validate this hypothesis by visualizing the value trade-off patterns of the top redditors.
[…]
The heat map tells us that there are some values commonly cherished by most of the redditors, yet they show different priorities and trade-offs for other values. For instance, when looking at the third column (tradeoff_C), almost all redditors prioritize “Effective Communication and Relationships” over “Emotional Influence and Accountability”. Redditors show various patterns when it comes to “Personal Autonomy and Freedom” (tradeoff_F to tradeoff_I), and especially for redditor 8, this value is prioritized in most cases, unless it conflicts with “Respect and Compliance with Authority”. This visualization not only suggests that active redditors in r/AmITheAsshole show distinct subjectivity, but explains their value preferences and judgments.
Figure 3: Euclidean distance of pairs of value representations with respect to their corresponding Pearson correlation coefficient computed with value win rates.
6.4 Do people show consistency in value preferences?
The idea of considering past comments as subjective ground and using them to infer redditors’ likely judgments on unseen situations is logically valid only if redditors exhibit consistent judgment patterns. For instance, if a redditor consistently favors one value, A, over a conflicting value, B, it is reasonable to infer that they would also prefer a similar value, A, assuming their decision-making remains consistent.
To show the consistency of value preferences, we plot the distance between pairs of value representations with respect to their corresponding correlation coefficients. We compute Pearson correlation coefficient using value win rates; when a value has a high win rate, another value that has correlation coefficients close to 1 also has a high win rate. Thus, more similar values should have correlation coefficients near 1, while increasing distance between values is expected to result in coefficients approaching -1.
Figure 3 illustrates the pairwise Euclidean distance of value representations on the x-axis and their Pearson correlation coefficients on the y-axis. The regression plot indicates a clear trend: as the similarity between two values increases, their correlation becomes more positive, whereas further distances correspond to more negative correlations. This analysis suggests that redditors show consistent value preferences, thereby validating our approach of using one’s past comments to infer their likely judgments.
7 Related Studies
The closest neighbor of this research is individual subjectivity analysis. Lee and Goldwasser (2022) analyzes the same community, r/AmITheAsshole and learned subjective preference of individuals by computing attention weights between rules-of-thumb and situations. In the political framing and agenda-setting domain, Roy and Goldwasser (2021) utilized Moral Foundations theory to characterize real world politicians varying in topics. Representing individual-level perspectives is also done in debate setting; Li et al. (2018) used graph embeddings to associate opnions and individuals together.
More recently, researchers utilized LLMs to ground their generations to a desired persona or personality. This persona-based control is done in many domains including basic question answering (Zheng et al., 2024a), interactive simulacra (Park et al., 2023), and solving causal inference and moral dilemma (Nie et al., 2023). Choi and Li (2024) proposed Persona In-Context Learning which uses Bayesian inference to select the optimal set of persona for a given task. Researchers have also tried providing more direct signals in the prompt using demographic information (Durmus et al., 2023).
Another line of related studies is the Retrieval-Augmented Generation. RAG retrieves the most relevant information to mainly fine-tune language models or help off-the-shelf LLMs infer better on many downstream tasks such as QA (Shi et al., 2023; Xu et al., 2023; Wang et al., 2024b), reasoning and language understanding (Yu et al., 2023; Lin et al., 2023; Zhang et al., 2023), text summarization and generation (Guo et al., 2023; Jiang et al., 2023; Yan et al., 2024), etc. Researchers have also studied whether RAG can be applied to more personal use case such as recommendation system (Rajput et al., 2023) and personalized dialog generation (Wang et al., 2023, 2024a). These approaches only consider factual information (e.g. purchase history, where did the person go two days ago) as a personalized aspect, while our research characterizes subjective perspectives and applies RAG for performance improvement.
8 Conclusion
In this paper, we propose a framework, SolAr, that takes redditors’ past comments into account and characterizes their subjective ground using value abstraction. Empirical results show that LLMs can be efficiently used to account for subjective preference of individuals compared to traditional methods that require fine-tuning. We also show that the performance can be further improved by retrieving the most relevant subjective ground using value trade-offs. Furthermore, SolAr provides an additional explanations about each redditor’s distinct value preference patterns which could later be used to justify LLM inference.
Limitations
Although it is not feasible to model subjectivity of individuals that is perfectly correct and inclusive, representing it with their past comments is still an over-simplified definition of subjectivity. In the future studies, we plan to incorporate more redditor-related information such as their community membership (i.e., what other subreddits they are actively participating) and activities on other communities to better characterize individuals.
Another limitation of this study is that the datasets and the tasks are tested only on a specific subreddit. Although r/AmITheAsshole is a huge online community that covers a wide range of situations exhibiting diverse perspectives, the usefulness of our framework will be more strongly validated when we apply our approach to other communities that require subjective perspectives.
[…]
Lastly, more validations of the generated abstract values are needed. As we use LLMs to freely generate value trade-offs of redditors that are observed from their comments, evaluating the soundness and quality of the values would make the subjectivity representation more powerful and useful. We further plan to validate this process with actual human evaluations and see if LLM-generated values can characterize human values reasonably well. The best way to evaluate LLMs’ characterization of individuals would be directly ask it to the target individuals. We leave recruiting human participants to accurately evaluate LLM generations as future work.
Ethics Statement
To the best of our knowledge, this work has not violated any code of ethics. All redditor information is anonymized in this paper as well as in the datasets we share to the public. The redditors are selected purely based on how active they participate in the community thus there is no discrimination in choosing redditors of interest. We denote that this paper poses potential risks where LLMs could misrepresent the subjectivity of individuals by referring to a limited number of past history and making hypotheses on their likely reactions to an unseen situations. We provide the code and datasets for future reproduction.
Appendix A Value Abstraction
A.1 Schwartz’s Theory of Basic Human Values
Schwartz (1992) defines theory of basic human values as:
  • Self-direction: “independent thought and action—choosing, creating, and exploring”
  • Stimulation: “excitement, novelty and challenge in life”
  • Hedonism: “pleasure or sensuous gratification for oneself”
  • Achievement: “personal success through demonstrating competence according to social standards”
  • Power: “social status and prestige, control or dominance over people and resources”
  • Security: “safety, harmony, and stability of society, of relationships, and of self”
  • Conformity: “restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms”
  • Tradition: “respect, commitment, and acceptance of the customs and ideas that one’s culture or religion provides”
  • Benevolence: “preserving and enhancing the welfare of those with whom one is in frequent personal contact (the ’in-group’)”
  • Universalism: “understanding, appreciation, tolerance, and protection for the welfare of all people and for nature”
In a more generalized view of the values, these values can be grouped into four categories, Openness to Change (self-direction, stimulation), Self-Enhancement (hedonism, achievement, power), Conservation (security, conformity, tradition), and Self-Transcendence (benevolence, universialism). Openness to Change and Conservation contradicts to each other, and Self-Enhancement and Self-Transcendence contradicts to each other as well.
In order to annotate comments to these fixed values, we first prompt gpt-4o model to generate values observed from the comments in general. Then we prompt again, asking “map the value items to the most relevant dimensions that Schwartz has defined in his Theory of Basic Human Values”.
A.2 Clustering Value Conflicts and Trade-offs
Figure 4: Visualization of Umap embeddings of value representations.
We follow the approach used in Lam et al. (2024) to cluster the value representations.
For the generated values that are conflicting in the situations in the dataset, we first obtain their vector representations. We use Open AI’s text embedding 3 small model (OpenAI, 2024), and reduced the dimensions to 256. We then further reduce the dimensionality using umap embeddings (McInnes et al., 2018). Figure 4 shows the 2-d visualization of umap embedding of all value representations, with number of neighbors as 30 and minimum distance as 0.
After this step, we use HDBSCAN to generate initial clusters while enforcing minimum of 100 items in each cluster. After obtaining the initial cluster, we gather all uncategorized values. We then compute the distance between each of the uncategorized value and initial cluster representations. If the distance is closer than a threshold (0.95), we assign the item to the cluster. For values that are not close enough to any other clusters, we group them using Kmeans clustering.
After assigning all value representations to a corresponding cluster, we ask gpt-4o-mini model to come up with a summary of unifying themes and patterns. Below is the template for the prompt:
I have this set of bullet point summaries of text examples: {examples} Please write a summary of unifying patterns for these examples. For each high-level pattern, write a 5 word NAME for the pattern and an associated one-sentence ChatGPT PROMPT that could take in a new text example and determine whether the relevant pattern applies. Please also include 2 example_ids for items that BEST exemplify the pattern. Please respond ONLY with a valid JSON in the following format : {  "patterns": [   {    "name": ,    "prompt": ,    "example_ids": ,   }   {    "name": ,    "prompt": ,    "example_ids": ,   }  ] }
Appendix B Model Implementation
B.1 Hyperparameter Searching
For all fine-tuned language models, we perform hyperparameter searching on training batch size and learning rates, and the best working combination is 16 and 1.895e-5. When adopting LoRA, we also explored r and alpha values, and we find that $r=4$, $alpha=16$ works the best.
B.2 Computing Resources
For encoder-only models, fine-tuning 8 models for each redditor takes approximately 90 to 160 minutes on a single Tesla V100 GPU. Fine-tuning encoder-decoder models, BART and FLAN-T5, took around 4 hours and 6 hours, respectively.
For inference using gpt-4o models, each of the experiment costs around USD 20, where the costs for input query takes about USD 19.90 and the rest is for the output which is either 0 or 1.